### v1

# Pressure-driven Valence evolution Coupled Hardening-to-Softening transition in YbPd

Abstract
We investigate the Yb valence instabilities in the strongly correlated YbPd compound using resonant X-ray emission spectroscopy under pressure across the charge-ordering (CO) transition. At a low temperature (T = 30 K) in the CO ordered phase, the Yb $4f$ valence remains nearly constant up to a pressure PK = 1.5 GPa, and then increases gradually at higher pressures. In contrast, at room temperature in the normal phase, an anomalous decrease of the Yb $4f$ valence is observed, without any accompanying structural phase transition. This behavior is corroborated by a systematic pressure dependent decrease of the unit-cell volume. Based on a Birch-Murnaghan analysis, the compressibility indicates hardening of the lattice with applied pressure up to a distinct kink seen at PK = 1.5 GPa. In contrast, for P >> PK, the Yb $4f$ valency saturates and the compressibility reveals a counterintuitive pressure-induced softening. The results show a minimum in the compressibility of YbPd (with $f^{13}$-$f^{14}$ hole-type mixed-valence) and is reminiscent of the maximum in compressibility seen in the $\gamma$-$\alpha$ first-order isostructural phase transition in cerium (with $f^{0}$-$f^{1}$ electron-type mixed-valence).
I INTRODUCTION
Cerium and ytterbium-based Kondo-lattice systems have attracted considerable attention over the past two decades owing to their outstanding physical properties, including unconventional superconductivity often associated with quantum phase transitions[39, 25]. These original properties are due to the competition between the Ruderman-Kittel-Kasuya-Yosida (RKKY) interaction and the Kondo effect. These properties may be tuned through control parameters such as chemical doping, pressure, or magnetic field, and the competition between the two interactions is well summarized in the Doniach phase diagram [15]. When the hybridization between the 4f states and the conduction band is weak, the RKKY interaction dominates. This leads to a stable trivalent 4f state and to the formation of large local magnetic moments, which order magnetically at low temperatures. As the hybridization strength increases, the Kondo interaction becomes dominant: the 4f magnetic moments are screened by the conduction electrons, and the valence deviates from the trivalent state.
The strongly correlated YbPd material is a unique Yb compound that exhibits both intermediate valence behavior and charge ordering (CO). It crystallizes in the cubic CsCl structure at room temperature and exhibits an antiferromagnetic phase transition at $T_{N}$ = 1.9 K due to the competition between the RKKY interaction and the Kondo effect [45]. It undergoes an incommensurate CO at $T_{1}$ = 130 K. At $T_{2}$ = 105 K, the incommensurate CO turns into a commensurate CO. With decreasing temperature, the cubic structure with a unique Yb site transforms into a tetragonal symmetry[8, 28] at 130 K, characterized by two distinct Yb sites. Below 105 K, a trivalent magnetic Yb3+ and strongly intermediate valent nonmagnetic Yb2.6+ are arranged alternatively along the $c$ axis, which has also been termed valence order[48, 51, 49, 47]. These findings are summarized in Fig. 1. The valence order in YbPd is very intriguing since YbPd shows metallic electrical resistivity from above $T_{1}$ and down to low temperatures. Hence, many studies have been devoted to the manifestation of the magnetic ordering in the intermediate valence YbPd compound and to the origin of the CO transition[45, 49, 1, 43, 2, 53]. More recently, the Kondo effect has been proposed as a mechanism linking the structural transition, the mixed valency, and zero thermal expansion in the CO phase[36].
Pressure-dependent ac-calorimetry and electrical resistivity measurements revealed another peculiar behavior of the YbPd compound. Unlike typical ytterbium compounds, where pressure usually stabilizes magnetic ordering, in YbPd the magnetic transition temperature $T_{N}$ decreases with increasing pressure and is discontinuously suppressed at PC = 1.9 GPa [48, 43, 2]. The CO transition at $T_{2}$ decreases monotonically with increasing pressure and also disappears at PC. A precise understanding of the exotic properties of YbPd requires detailed knowledge of the temperature and pressure dependence of the Yb valence. While temperature-induced valence changes have been investigated using high-energy spectroscopic techniques, the pressure dependence of the valence in YbPd has not been reported.
In this study, we investigated the pressure dependence of the Yb valence in YbPd under high pressure at temperatures below and above the CO transition by means of High-resolution Fluorescence detected X-ray absorption spectroscopy (HERFD-XAS) and Resonant X-ray Emission Spectroscopy (RXES) at the Yb L3 edge which are sensitive probes of 4f valence state owing to their superior spectral resolution[12, 31, 32]. To study the interplay between the Yb valence and the CO, we measured the pressure (P) dependence of the Yb valence across the CO transition (at 30 K), and in the cubic phase (at room temperature). At low temperature (T = 30 K), the valence is constant up to 1.5 GPa, and it increases with P in the cubic phase. Remarkably, the measurements at room temperature revealed a significantly large decrease of the Yb valence at low pressures P<< PK, followed by a nearly constant mixed valency at higher P in the cubic phase. This is corroborated by a kink observed at P = PK in the P-dependent unit cell volume. Moreover, whereas the compressibility of most materials generally increases with P, YbPd exhibits a P-dependent minimum in its compressibility at PK. This behavior is reminiscent of the $\gamma$-$\alpha$ first-order isostructural phase transition in cerium metal , albeit, the $\gamma$-$\alpha$ transition exhibits a maximum in its compressibility.
II EXPERIMENTAL DETAILS
The polycrystalline YbPd sample was prepared from high purity (at least 99.9 %) commercially available elements. A mixture with Yb1.03Pd starting composition to compensate for the Yb vaporization at elevated temperature was placed in a Ta crucible. The crucible was sealed by arc-welding in a glove-box under pure Ar atmosphere. The crucible was heated up to $\sim$ 1550°C in an induction furnace for several minutes. Room-temperature x-ray diffraction indicates that the sample is single-phase with the CsCl type cubic structure. The RXES and HERFD-XAS measurements were performed using the multi-analyzer x-ray emission spectrometer ”MULTIXS” on the RIXS endstation of the GALAXIES beamline[4] at the SOLEIL synchrotron. The spectrometer was equipped with three Si(620) bend crystal analyzers (with a bending radius R = 1 m) which were operated around the Bragg angle of 76.770°. The overall energy resolution was about 1 eV as determined by the full width at half maximum (FWHM) of the elastic line. To study the P-T dependences of the Yb valence, the sample was mounted in a diamond anvil cell with silicon oil as a pressure-transmitting medium and ruby as a pressure indicator. The cell was subsequently mounted in a specially designed high-pressure He flow cryostat.
III RESULTS
Figure 1: Crystal structure of YbPd at ambient pressure above (a) $T_{1}$ = 130 K (cubic phase) and below $T_{1}$ (tetragonal phase). YbPd shows only one-type of Yb-site in the normal cubic phase, but shows 2 types of Yb-sites in the tetragonal CO phase. P-T phase diagram reported in Ref [43].
Fig. 2a illustrates the pressure-dependent Yb-L3 HERFD-XAS spectra of YbPd measured at $T=30$ K. The spectra were recorded by scanning the incident photon energy through the Yb L3 absorption edge while monitoring the scattered intensity of the Yb Lα1 fluorescence line ($2p^{5}3d^{10}\rightarrow 2p^{6}3d^{9},h\nu_{\text{out}}=7417\,\text{eV}$). The HERFD-XAS spectral shape is sharper than the standard XAS spectrum because the broadening is determined by the longer lifetime of the shallower 3dd hole, rather than by that of the deep 2pp hole[27]. The spectra exhibit two peaks at around 8940 eV and 8947 eV corresponding to the transition from the initial intermediate valent ground state $\big|g\big\rangle=a\big|f^{13}\big\rangle+b\big|f^{14}\big\rangle$ to a final state with mainly Yb2+ ($2p^{6}4f^{14}\rightarrow 2p^{5}4f^{14}5d^{1}$) and Yb3+ ($2p^{6}4f^{13}\rightarrow 2p^{5}4f^{13}5d^{1}$) characters. This intermediate valence character at low temperatures is consistent with early XAS at Yb $L_{3}$ edge (average valence $v=$ 2.80) [45] as well as soft x-ray photoemission spectroscopy measurements ($v=$ 2.75)[14], while hard x-ray photoemission spectroscopy measurements showed $v=$ 2.79[36]. The Yb2+ structure is split into two peaks separated in energy by $\sim$2.3 eV. This splitting is attributed to the 5d crystal field effect[55, 30]. With increasing pressure, a clear spectral-weight transfer from Yb2+ to Yb3+ is observed. However, at room temperature, the Yb3+ decreases and is transferred to Yb2+ with increasing pressure as shown in Fig. 2b.
Figure 2: Yb L3 XAS and RXES spectra of YbPd as a function of temperature at (a) $T$ = 30 K and (b) $T$ = 300 K.
Figure 3: Yb L3 RXES spectra of YbPd as a function of temperature at (a) $T$ = 30 K and (b) $T$ = 300 K. The incident photon energy was set to the maximum of the Yb2+ resonance at the Yb L3 absorption edge and the spectra were normalized to the areaa under the curve.
Figure 4: (a) Evolution of the Yb valence as a function of pressure at T = 30 K. Evolution of the Yb valence (b), the compressibility (b) and the Kondo temperature (c) as a function of pressure at room temperature.
[…]
$v=3-\frac{I^{(2+)}}{I^{(2+)}+\frac{I^{(3+)}}{\zeta}},$
where $I^{(2+)}$ and $I^{(3+)}$ are, respectively, the integrated intensities of the peaks Yb2+ and Yb3+.The parameter $\zeta$ accounts for the reduced Yb3+ spectral weight observed in RXES spectra, which depends on the incident photon energy. A value of $\zeta=8.5\%$ yields a valence at ambient pressure consistent with that determined by hard X-ray photoemission spectroscopy [36].
Fig. 4a shows the pressure dependence of the Yb valence obtained from RXES at room temperature and at $T=$ 30 K. Interestingly, two pressure ranges can be identified: below 1.5 GPa, the valence shows a negligible change, while above 1.5 GPa, the valence increases with increasing pressure. Earlier work [43] showed that for a pressure of about 1.5 GPa, the commensurate CO sets in at $T_{2}$ = 30 K (Fig. 1(c)). More importantly, at room temperature, we found an anomalous pressure-induced change in the Yb valence. The pressure dependence of the Yb valence is characterized by a sharp initial drop at low pressures (P $\leq$ 1.5 GPa), where the valence rapidly decreases from 2.87 to 2.58 and remains constant with a further increase in pressure. In typical Yb compounds, the valence usually increases with increasing pressure since the ionic radius of the Yb3+ ($f^{13}$) ion is smaller compared to that of the Yb2+ ($f^{14}$) ion[11, 22, 46, 40, 42, 18, 17, 50, 23, 3]. A similar pressure-induced anomalous valence reduction on increasing pressure has recently been observed in YbCu5 [57], YbCu4.5[58], and YbCu6.5[56].
It is well known that, in Yb-based compounds, the Kondo temperature $T_{K}$ scales with the occupation of the $4f$ hole number $n_{h}$ (the Yb valence $v=2+n_{h}$). By combining the Kondo energy scale $T_{K}$, extracted from thermodynamic measurements, with the $4f$-state occupation determined by RIXS experiments, Kummer et al. [32] recently showed that this scaling follows a universal power law, $3-v=aT_{K}^{n},$ with $n=2/3$ and $a=1/200$. We employed this power law to estimate the pressure dependence of the Kondo temperature for the single Yb site at room temperature, as shown in Fig. 4c. The Kondo temperature increases markedly from 137 K to 768 K at $P_{K}=1.5$ GPa and then remains nearly unchanged upon further increase in pressure.
IV Discussion
The pressure dependence of the Kondo temperature TK in Ce- and Yb-based heavy-fermion compounds has been studied theoretically within the framework of the single impurity Anderson model (SIAM) [26]. The calculations showed that for Ce compounds, TK increases monotonically, while for Yb compounds, it decreases at low pressure, shows a broad minimum, and then increases with further increasing pressure. YbPd may be located near the pressure of the broad minimum. Indeed, at ambient pressure, the valence of the Yb1 site initially increases and becomes magnetic and trivalent as a result of the pressure induced by the displacement of the Pd atom. Moreover, the magnetic transition temperature $T_{N}$ decreases under pressure. Therefore, applying pressure results in transforming the Yb1 site from a nearly trivalent state towards an intermediate valence state. Simultaneously, the pressure independence of the Yb valence in the CO phase suggests that the decrease in the Yb1 site valence is compensated by a corresponding increase at the Yb2 site valence.
The anomalous valence evolution under pressure in YbPd is corroborated by the pressure dependence of the unit cell volume. Fig. 5a shows the contour map of X-ray diffraction intensities as a function of pressure. Fig. 5c shows the pressure dependence of the unit cell volume extracted from the XRD data. The unit cell volume decreases with pressure, and a kink is observed at around 1.5 GPa. Indeed, as pressure is applied, the Yb ions in YbPd transform from Yb3+ to Yb2+, which brings about a local volume expansion at the same time as the pressure is trying to reduce the unit cell volume. This competition results in a hardening of the lattice because of the increase in the Yb2+ fraction, as we show below from a Birch-Murnaghan analysis. On further increase in the applied pressure, at about 1.5 GPa, the valence saturates, and hence the volume change due to valency induced expansion disappears, and the rate of volume reduction incrases (Fig. 5c). This results in the observed kink at P = PK, indicating a change in stiffness which we quantify using a Birch-Murnaghan analysis in the following.
Figure 5: (a) Contour map of X-ray diffraction intensities collected in the pressure range 0–3 GPa. (b) XRD patterns at P = 0.2 GPa and 2 GPa. (c) Pressure dependence of the volume. The blue and red curves are the to third-order Birch- Murnahgan’s equation of state for P $\leq$ PK and P $\geq$ PK respecively. The arrow indicates the presure PK where the kink is observed.
We extracted the Bulk modulus $K_{0}$, which characterizes the compression behavior under hydrostatic pressure, and its pressure derivative $K^{{}^{\prime}}_{0}=\frac{\partial K_{0}}{\partial p}$ that defines the change in stiffness using the standard third-order Birch-Murnaghan equation of state [7]:
$P(V)=\frac{3}{2}K_{0}$ $\left[\left(\frac{V_{0}}{V}\right)^{\frac{7}{3}}-\left(\frac{V_{0}}{V}\right)^{\frac{5}{3}}\right]$ (1)
$\times\left[1+\frac{3}{4}(K_{0}^{\prime}-4)\left(\left(\frac{V_{0}}{V}\right)^{\frac{2}{3}}-1\right)\right]$
where P and V are the pressure and volume, respectively, and V0 and K0 are the values of V and K at zero pressure. The data for the V(P) functions, illustrating the fitted equation of state, are shown in Figure 5c. We obtain the parameter of $K^{{}^{\prime}}_{0}$ = 22.3, $K_{0}$ = 47.18 GPa, $V_{0}$ = 40.66 Å3 for P << PK and $K^{{}^{\prime}}_{0}$ = -6.1, $K_{0}$ = 38.32 GPa, $V_{0}$ = 39.6 Å3 for P >> PK. Remarkably, we observe a sign change in the derivative of the bulk modulus $K^{{}^{\prime}}_{0}$, which is a signature of a first-order isostructural transition [33]. Indeed, for P << PK, the compressibility $\beta=-\frac{1}{V}\frac{\partial V}{\partial p}$ decreases with increasing pressure as is typical for solids [35], while for P >> PK, the compressibility follows an unexpected monotonic increase with pressure as shown in Figure 4b. Negative values of $K^{\prime}_{0}$ are very rare and correspond to an effect known as pressure-induced softening [52, 16, 44, 20, 21, 19, 10, 54], a counterintuitive phenomenon in which a material becomes increasingly compressible under pressure.
Pressure-induced softening has been observed in a few materials, including cerium metal[9], the amorphous silica [52], ZrW2O8[44], Zn(CN)2[20] and ScF3[54]. Among these materials, cerium exhibits strong similarities with YbPd. Indeed, cerium undergoes an isostructural face-centered cubic (fcc) $\gamma$-$\alpha$ phase transition, accompanied by a volume collapse of approximately 15% at room temperature under pressure. The compressibility of cerium exhibits a maximum at the $\gamma$-$\alpha$ volume-collapse transition; it increases with pressure in the $\gamma$-phase and decreases in the $\alpha$-phase [9, 29, 37, 13, 38]. Moreover, the phase transition in Ce is marked by a pronounced increase in the Kondo temperature, occurring concomitantly with the volume collapse under pressure. The $f$ electrons play a central role for this transition. Both the pressure-temperature phase diagram and the temperature dependence of the bulk modulus of cerium have been succesfully repoduced within the Kondo volume collapse model [6, 34, 5] in which a rapid pressure-induced increase of the Kondo temperature drives the transition. As shown in Fig. 4c, YbPd likewise exhibits a rapid increase in the Kondo temperature at low pressures upto PK at which the compressibility reaches a minimum, and then becomes nearly constant above PK. A theoretical analysis based on the Kondo volume collapse model is essential to test its applicability in describing the pressure-induced changes of the strongly correlated YbPd compound.
V CONCLUSION
We successfully investigated the pressure dependence of the Yb valence in YbPd using Yb-L3 resonant inelastic X-ray emission spectroscopy. Our RXES data at room temperature showed that the Yb valence decreases from 2.87 to 2.58 at 1.5 GPa and subsequently saturates with further increase of the pressure. The results reveal a pressure-driven valence evolution which is coupled to an unexpected hardening-to-softening transition in YbPd as a function of pressure. We evidenced a pronounced minimum in the pressure dependence of the compressibility suggesting a first order isostructural transition.
VI ACKNOWLEDGMENTS
We acknowledge the French Synchrotron facility SOLEIL (Saint-Aubin, France) for the allocated beam time (exp. n° 20241507).